A number is 64 times the square of its reciprocal. What is the number?
Answer by Theo(13342)

let the number = x.

x = 64 * (1/x)^2.

this is equivalent to x = 64 * 1^2 / x^2 which is equivalent to x = 64 * 1/x^2.

multiply both sides of the equation by x^2 to get x^3 = 64

take the cube root of both sides of the equation to get x = 4.

when x = 4, the equation of x = 64 * (1/x)^2 becomes 4 = 64 * (1/4)^2 which becomes 4 = 64 * 1/16 which becomes 4 = 4, confirming the solution is correct.